Joyce Corporation acquired 60 percent ownership of Ria Company on January 1, 2014 for P525,000. The purchase differential was assignable entirely to a depreciable asset with a remaining economic life of 10 years at the date of acquisition. On that date, the following amounts were reported in the shareholders’ equity by Ria Company

                        Ordinary Shares                                              P300,000

                        Capital in Excess of par value                         150,000

                        Accumulated profits                                         300,000

In the consolidated statement of financial position prepared as of December 31, 2017, the non-controlling interest was reported at P348,000. Joyce reported operating income of P150,000 during the 4-year period and paid dividends of P67,500. Joyce had accumulated profits of P900,000 on the date of combination.

Required:

Determine the consolidated accumulated profits as of December 31, 2017.
Answer #1

Joyce Corp. = Parent Company, holding 60% of Ria Co.

Ria Co. = Subsidiary

Consolidated Accumulated Profits as on 31.12.17:

Accumulated Profits of Joyce as on 31.12.17= P900000+P150000-P67500=P982500

Non Controlling Interest on 31.12.17=40%= P348000

So, Holding 60%=348000/40%*60%= P522000

Capital in Excess of Par Value = P150000

Acc. Profit of Ria as on acquisition = P300000

Thus Consolidated Acc. Profits = 982500+522000+150000+300000

= P1954500